OpenStudy (anonymous):

((2^6) x (2^4)) / (2^2)

OpenStudy (shamil98):

\[\huge 2^6 * \frac{ 2^4 }{ 2^2 } = 2^6 * 2^2 = 2^8\]

OpenStudy (anonymous):

Thanks

OpenStudy (shamil98):

Look at this for explanation on the rules: http://www.purplemath.com/modules/exponent.htm